Last year,tom was thrice as old as his sister.next year,he will be twice as old as his sister.what will be tom's age when his sister is 12 years old?
Answer by ankor@dixie-net.com(22740)

Let t = Tom's present age
Let s = Sister's present age
:
Last year,tom was thrice as old as his sister.
t-1 = 3(s-1)
t-1 = 3s - 3
t = 3s - 3 + 1
t = 3s - 2
:
next year,he will be twice as old as his sister.
t + 1 = 2(s+1)
t + 1 = 2s + 2
t = 2s + 2 - 1
t = 2s + 1
replace t with (3s-2)
3s - 2 = 2s + 1
3s - 2s = 1 + 2
s = 3 yrs is sisters age
then
t = 3(3) - 2
t = 7 yrs is Tom's age
:
what will be tom's age when his sister is 12 years old?
Sister will be 12 in 9 years
therefore
7 + 9 = 16 yrs will be Tom's age